OpenStudy (anonymous):

HELP PLEASE! SHOW ME HOW TO DO IT AND NOT JUST GIVE ME THE ANSWER. 27x4 – 90x2 + 75

OpenStudy (anonymous):

might the question be "factor \(27x^4 – 90x^2 + 75\)"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factor out a 3 from the whole thing first

OpenStudy (anonymous):

then we can proceed, assuming you don't want just the answer

OpenStudy (anonymous):

9 - 30+25

OpenStudy (anonymous):

don't forget the variables and the 3 out front! \[3(9x^4-30x^2+25)\]

OpenStudy (anonymous):

I didnt, i was just giving you the numbers

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

ok now you are supposed to recognize the second part as a perfect square \[(a-b)^2=a^2-2ab+b^2\] and in this case \(a=3x^2,b=5\)

OpenStudy (anonymous):

so no, that is not the final answer, that is the first step

OpenStudy (anonymous):

final answer is what you get when you make the above replacements, namely \[3(3x^2-5)^2\]

OpenStudy (anonymous):

@satellite73 can you help me with 2 more please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

32x6 – 162

OpenStudy (anonymous):

\[32x^6-162\] common factor of \(2\) gives \[2(16x^6-81)\]

OpenStudy (anonymous):

this time you are supposed to recognize \(16x^6-81\) as the difference of two squares \[a^2-b^2=(a+b)(a-b)\] with \(a=4x^3, b=9\)

OpenStudy (anonymous):

final answer here is therefore \[2(4x^3+9)(4x^3-9)\]

OpenStudy (anonymous):

clear enough?

OpenStudy (anonymous):

yes. TY. one more lol

OpenStudy (anonymous):

5x3 – 35x2 + 60x

OpenStudy (anonymous):

this time factor out the common factor of \(5x\)

OpenStudy (anonymous):

start with \[5x(x^2-7x+12)\]

OpenStudy (anonymous):

then the question is how to factor \(x^2-7x+12\) do you know how or would you like me to show you?

OpenStudy (anonymous):

(x )(x ) find the factors of 12 that add up to -7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

obviously they are both negative, since \(-7\) is negative and \(+12\) is positive

OpenStudy (anonymous):

(x-4)(x-3)

OpenStudy (anonymous):

you got it final answer is \[5x(x-4)(x-3)\]